OpenStudy (luigi0210):

Differentiate: \[\LARGE (1-4x+7x^5)^{30}\]

OpenStudy (anonymous):

by chain rule set \[u=1-4x+7x^5,\frac{du}{dx}=-4+35x^4\] \[y=(1-4x+7x^5)^{30}=u^{30},\frac{dy}{du}=30u^{29}\]and use \[\frac{dy}{dx}=\frac{du}{dx}\frac{dy}{du}\] and substitute

OpenStudy (anonymous):

\[30(1-4x+7x5)^{29}(1-4x+7x^{30})'\]

OpenStudy (anonymous):

^^^ alternative method

OpenStudy (luigi0210):

Makes sense, thanks :)

OpenStudy (anonymous):

@Luigi210 and @Luigi0210 Can you help me in health scicene

OpenStudy (anonymous):

O.O

OpenStudy (luigi0210):

Another me? Great.

OpenStudy (anonymous):

you have two accounts??? lol the one is 16 years old the other is 99

OpenStudy (anonymous):

Can you?

OpenStudy (anonymous):

um i guess

OpenStudy (anonymous):

Okay go to heatlh scicens

OpenStudy (anonymous):

theres a typo in the derivative \[30(1-4x+7x^5)^{29}((1-4x+7x^5)^{30})'\]

OpenStudy (owlcoffee):

It's a composite function and we want it's derivative: Let's begin by recognizing the functions: \[(1-4x+7x ^{5})^{30}\] I said in an other question, that the derivative of a composite function is: \[y=g(f(x))^{m}\] \[y'=m.g(f(x))^{m-1}.f'(x)\] meaning that the derivative of the function you showed is: \[y'=30(1-4x+7x ^{5})^{29}(35x ^{4}-4)\]